When 0.08 moles of NaOH into 500 ml of 0.6 M buffer system at pH 6.6, the ratio of [Salt]/[Acid] becomes 0.1 (adding NaOH does not change volume). What is the pKa of the buffer system?
1
Expert's answer
2020-11-23T06:53:07-0500

Amount of buffer = 0.6M × 500/1000L = 0.3 mol


When 0.08 moles of NaOH reacts with 0.3 moles of buffer (let's assume the buffer is mono protic)


NaOH is the limiting reagent and the amount of buffer left is 0.3 - 0.08 = 0.22 moles


Concentration of NaOH unreacted = 0.22/500ml = 0.44M


Comparing with the ratio above,

Salt/Acid = 0.1

Since salt = 0.08, acid = 0.8 moles = 0.88/500ml = 1.76M


from,

pKa = pH - log"\frac{[Acid]}{[Conjugate\ base]}"


pKa = 6.6 - log"\frac{1.76}{0.44}"

pKa = 6.6 - 0.602 = 5.998